15, O -1 points Tiplere 1. F.070. If the average d
Answer #1

11.

a) (1.16)(9.99x10^4) = 11.5884x10^4

b) (2.86x10^-8 ) - (5.31x10^-9) = 2.329x10^-8

c) 12pi/ (4.76x10^-3) = 7919.98

d) 24.8 + (5.99x10^2) = 623.8

9.

the momentum of the object is the product of mass and velocity.

now momentum = mass x velocity = mass x (velocity/time) x time

= mass x acceleration x time = force x time